# Coarse-to-Fine Personalized LLM Impressions for Streamlined Radiology Reports

### v1

@@ 1 Introduction
This research explores the application of large language models (LLMs)  (Li et al., 2025; Cai et al., 2024; Yao et al., 2025) to a broader range of domains. Specifically, we investigate their use in summarizing radiology reports, addressing critical challenges in clinical decision-making and workflow efficiency. Artificial Intelligence (AI) and Natural Language Processing (NLP) have become integral to healthcare, enabling advancements in patient care, clinical workflows, and medical research. Applications include automating clinical documentation, such as extracting information from clinical notes or summarizing patient histories, as demonstrated in this research. These tools also enhance clinical decision support systems, improving the overall clinical experience.
@@ 2.3 LLM and framework in healthcare
In the healthcare domain, large language models have first been fine-tuned to generate structured SOAP notes, as demonstrated by Leong et al.’s efficient fine-tuning approach (Leong et al., 2024b). Building on this work, MediNote is the first end-to-end generative AI framework that combines retrieval-augmented generation (RAG), parameter-efficient fine-tuning, and ambient listening to convert raw clinical conversations directly into SOAP documentation (Leong et al., 2024a). Complementing these advances, AgentNet (Leong & Wu, 2024) and (Wang et al., 2025) employs a fully decentralized, retrieval-augmented architecture in which autonomous agents dynamically reconfigure their roles and interconnections within a directed acyclic graph to optimize clinical workflows and improve patient outcomes in healthcare. Together, these developments chart a clear progression—from LLM fine-tuning for note generation, through RAG-enhanced generative frameworks, to autonomous multi-agent systems—highlighting the evolving landscape of AI-driven healthcare solutions.
@@ 4.4 Coarse-to-Fine Generation Process
The Coarse Stage initiates the summarization process by utilizing large language models (LLMs) to generate preliminary impressions based on the input findings and patient data. This stage is designed to capture primary clinical insights while ensuring the output follows a structured format. We designed a prompt engineering strategy that evolves from a simple base prompt to more detailed and customized versions tailored for specific audiences. The base model utilizes a straightforward prompt, such as "Summarize the following input"(Wang et al., 2023), which generates bullet-point key points. Building on this, we incorporated role descriptions and task-specific instructions to produce more detailed outputs while maintaining the bullet-point structure. Additionally, we provided three-shot examples to guide the model in generating outputs aligned with specific writing styles and standards. For expert-level outputs, the prompts further emphasize role descriptions, expanded task instructions, and a clear definition of the target audience to ensure the generated content meets their expectations. Users can also adjust the output length based on specific requirements. Full details of the prompts, including examples for base, detailed, and expert versions, are provided in Figure 3.

### v2

@@ 1 Introduction
This research explores the application of large language models (LLMs)  (Li et al., 2025b; Cai et al., 2024; Yao et al., 2025) to a broader range of domains. Specifically, we investigate their use in summarizing radiology reports, addressing critical challenges in clinical decision-making and workflow efficiency. Artificial Intelligence (AI) and Natural Language Processing (NLP) have become integral to healthcare, enabling advancements in patient care, clinical workflows, and medical research. Applications include automating clinical documentation, such as extracting information from clinical notes or summarizing patient histories, as demonstrated in this research. These tools also enhance clinical decision support systems, improving the overall clinical experience.
@@ 2.3 LLM and framework in healthcare
In the healthcare domain, large language models have first been fine-tuned to generate structured SOAP notes, as demonstrated by Leong et al.’s efficient fine-tuning approach (Leong et al., 2024b; Zhang et al., 2025a, b). Building on this work, MediNote is the first end-to-end generative AI framework that combines retrieval-augmented generation (RAG), parameter-efficient fine-tuning, and ambient listening to convert raw clinical conversations directly into SOAP documentation (Leong et al., 2024a; Wang et al., 2025a). Complementing these advances, AgentNet (Leong & Wu, 2024) and (wang2025one) employs a fully decentralized, retrieval-augmented architecture in which autonomous agents dynamically reconfigure their roles and interconnections within a directed acyclic graph to optimize clinical workflows and improve patient outcomes in healthcare. Together, these developments chart a clear progression—from LLM fine-tuning for note generation, through RAG-enhanced generative frameworks, to autonomous multi-agent systems—highlighting the evolving landscape of AI-driven healthcare solutions.
@@ 4.4 Coarse-to-Fine Generation Process
The Coarse Stage initiates the summarization process by utilizing large language models (LLMs) to generate preliminary impressions based on the input findings and patient data. This stage is designed to capture primary clinical insights while ensuring the output follows a structured format. We designed a prompt engineering strategy that evolves from a simple base prompt to more detailed and customized versions tailored for specific audiences. The base model utilizes a straightforward prompt, such as "Summarize the following input"(Wang et al., 2023a), which generates bullet-point key points. Building on this, we incorporated role descriptions and task-specific instructions to produce more detailed outputs while maintaining the bullet-point structure. Additionally, we provided three-shot examples to guide the model in generating outputs aligned with specific writing styles and standards. For expert-level outputs, the prompts further emphasize role descriptions, expanded task instructions, and a clear definition of the target audience to ensure the generated content meets their expectations. Users can also adjust the output length based on specific requirements. Full details of the prompts, including examples for base, detailed, and expert versions, are provided in Figure 3.